OpenStudy (anonymous):

divide the rational expression: x^2-9/5x+15 / 3-x/5

OpenStudy (anonymous):

x²-9/5x+15 / 3-x/5 = x²-9/5x+15 / 5/x-3 = (x+3)(x-3)/5(x+3) / 5/x+3 = x-3/3-x = -1